# backgroundcolor=, commentstyle=, breakatwhitespace=false, keywordstyle=, language=Java, stringstyle=, basicstyle=, showstringspaces=false

### v1

@@ Abstract
Technical debt (TD) is a term used to describe the additional work and costs that emerge when developers have opted for a quick and easy solution to a problem, rather than a more effective and well-designed, but time-consuming approach. Self-Admitted Technical Debts (SATDs) are a specific type of technical debts that developers intentionally document and acknowledge, typically via textual comments. While these self-admitted comments are a useful tool for identifying technical debts, most of the existing approaches focus on capturing crucial tokens associated with various categories of TD, neglecting the rich information embedded within the source code itself. Recent research has focused on detecting SATDs by analyzing comments embedded in source code, and there has been little work dealing with technical debts contained in the source code. To fill such a gap, in this study, through the analysis of comments and their associated source code from 974 Java projects hosted in the Stack corpus, we curated the first ever dataset of TD identified by code comments, coupled with its associated source code. Through an empirical evaluation, we found out that the comments of the resulting dataset help enhance the prediction performance of state-of-the-art SATD detection models. More importantly, including the classified source code significantly improves the accuracy in predicting various types of technical debt. In this respect, our work is two-fold: (i) We believe that our dataset will catalyze future work in the domain, inspiring various research issues related to the recognition of technical debt; (ii) The proposed classifiers may serve as baselines for other studies on the detection of TD by means of the curated dataset.
Index Terms:
Technical debt, mining software repositories, source code.
1 Introduction
@@
The concept of technical debt was originally introduced by Cunningham [15] to represent the liabilities that arise when developers make sub-optimal technical decisions, either intentionally or unintentionally during the software development life-cycle in their rush to market. Various factors can lead to the accumulation of technical debt, including deadline pressures, existing low-quality code, misaligned incentives, and poor software processes, among others [6]. Previous studies have shown that developers often underestimate the consequences of such debts, which can degrade the quality of the source code, increase bug rates, and slow development velocity [78, 72]. Identifying the code that contains technical debt is crucial to a rational development process, as this allows developers to fix the most important issues, the ones that are slowing the project down. Given the significant impact of technical debt on software maintenance and quality assurance, substantial research has been conducted to identify and analyze technical debt in software engineering [38, 21, 3, 20, 8].
@@
In 2015, da Silva Maldonado et al. [42] introduced their seminal work on detecting self-admitted technical debt (SATD) from comments embedded in source code. In particular, the authors manually classified a set of code comments to identify various types of technical debt. Afterward, they also developed various models to detect SATD from their dataset  [16]. Significant research on the recognition of SATD has flourished since then [18, 55]. Recently, there have been various approaches proposed to recognize technical debt contained in different project artifacts. Among others, Li et al. [37] conceived an approach to identify SATD from four independent sources, i.e., source code comments, commit messages, pull requests, and issue tracking systems. Tan et al. [64] manually curated a dataset of TD manifested in 3,000 issues. An evaluation of the collected dataset showed that there is a positive correlation between the number of TD items identified and mentioned as resolved in issue trackers and the number of debt items paid back in the source code.
@@
The majority of research thus far has mined TD from textual sources, e.g., comments [42], issues [64, 37], or pull requests [37]. In this respect, SATD detection tools rely heavily on text to function. If no technical debt is reported in text, a debt might exist, but these tools would fail to identify it. To add insult to injury, even when there are comments, many of them and the corresponding code are not coherent, i.e., comments may be out of date and incorrectly reflect what is actually contained in the associated code [14]. This may happen because developers forget to update their comments after they made changes to the code [28, 49]. To the best of our knowledge, there has been no attempt to create large-scale datasets of technical debt directly contained in source code. But we see a need for this type of data, to reduce the dependence on textual comments in detecting TD, thus vastly enhancing the contexts in which debt may be detected.
@@
To address these challenges, in this paper we propose a pipeline for the enrichment of technical debt data. In particular, our methodology involves extracting SATD comments in conjunction with corresponding source code units. We have devised a method to identify Java source code that possibly contains technical debt, creating our initial corpus. Then we manually classified five categories of technical debt in this corpus. In addition, we developed a machine learning based tool to detect technical debt contained in textual comments and source code. By means of an empirical evaluation, we demonstrated that the curated dataset has the potential to advance state-of-the-art research in the domain, paving the way for a completely new way of identifying TD.
@@
  • RQ1: Do the manually classified comments contribute to an improvement in the detection of SATD? We augmented the dataset collected by da Silva Maldonado et al. [42] with our newly classified comments to yield a combined dataset. Afterward, we ran four machine learning models, i.e., BERT, RoBERTa, UniXCoder, CodeBERT, on both datasets to compare the prediction performance of these models. This aims to determine whether the new comments are useful in predicting technical debt.
  • RQ2: Does the inclusion of source code help to enhance the detection of technical debt? We enriched the input data in RQ1 with source code and fine-tuned the four aforementioned machine learning models to identify technical debt in code. This RQ aims to investigate whether the enriched dataset (consisting of classified comments and corresponding source code) is beneficial to the detection of technical debt.
  • RQ3: What is the accuracy of different pre-trained models when detecting TD solely from source code? Using various machine learning models, we ran experiments on the collected dataset to investigate the accuracy of these models in classifying debt contained solely in source code. With this RQ we investigate to what extent existing deep learning algorithms are able to detect TD from code, thus inspiring future research in this direction.
@@ 2 Background
2.1 Self-admitted technical debt (SATD)
SATD is the technical debt that is expressly admitted by a developer through comments embedded in source code, issue trackers [36], commit messages, or pull requests [37]. da Silva Maldonado et al. [42] identified five types of SATD, i.e., DESIGN, DEFECT, DOCUMENTATION, REQUIREMENT/IMPLEMENTATION, and TESTING. This categorization allows for more insightful descriptions and a deeper understanding of the non-optimal solution options taken. This section takes various examples to explain these SATD categories.
@@ 2.1 Self-admitted technical debt (SATD)
2.2 Pretrained Language Models
Language models (LMs) are a foundational component in natural language processing (NLP) that have significantly advanced over the past decade. Recently, LMs have been powered by neural networks and trained on large text corpora, being able to capture both the syntactic and semantic aspects of languages more effectively. These models commonly follow the pre-training and fine-tuning paradigm [81]. During the pre-training phase, models are trained on large-scale unlabeled corpora using task-agnostic objectives such as word prediction, resulting in the development of pre-trained language models (PLMs). PLMs are then fine-tuned to adapt to various downstream tasks. Early PLMs [56] were mostly based on Recurrent Neural Networks (RNNs) and their variants, such as long short-term memory (LSTM) [33] and gated recurrent units (GRU) [10]. However, these approaches were computationally inefficient due to limitations in parallel processing, reducing scalability when training with extensive datasets and large model sizes. With the introduction of the Transformer architecture [68] and its self-attention mechanism, significantly more parallelization became possible as compared to RNNs. This advancement enables efficient pre-training of large language models on extensive datasets using multiple GPUs. Various transformer-based PLMs have achieved state-of-the-art performance across a wide range of tasks [17, 39, 31, 52, 65]. Given the superior performance of transformer-based PLMs, which have also been explored in the context of SATD detection [18, 58, 59], our study concentrates on utilizing these models.
2.2.1 Transformer architecture
This section provides an overview of the Transformer architecture, emphasizing key components and elements [68].
@@ 2.2.2 Types of PLMs
Based on the neural architectures of Transformer-based PLMs, we categorize the models into three main groups, as also outlined in existing work [44].
@@
Encoder-based PLMs: This type of model utilizes the Transformer Encoder and builds a network by stacking multiple layers. These models were initially developed for language understanding tasks, such as text classification, where the objective is to predict a class label for a given input text. The pre-training stage of these models typically involves corrupting a given sentence in some way (e.g., by masking random words) and then training the model to identify or reconstruct the original sentence. To tackle a downstream task such as sentence classification or named entity recognition, these models are fine-tuned on task-specific data, and this involves substituting the LM head (the word prediction layer), with a classification head. BERT [17], a prominent encoder-based model, has inspired the development of several variants, such as RoBERTa [39] and ALBERT [31], which have demonstrated substantial improvements across various understanding tasks.
@@
Bidirectional Encoder Representations from Transformers (BERT) [17] is among the most widely adopted encoder-based PLMs. During pretraining, BERT leverages two objectives: masked language modeling (MLM) and next sentence prediction (NSP). In MLM, random tokens within a sentence are masked, and the model is trained to predict these masked tokens using the context of the surrounding words. Meanwhile, NSP trains BERT to comprehend the relationship between two sentences by predicting whether one sentence logically follows the other. RoBERTa [39] extends BERT by improving its robustness through refined model design choices and training strategies. These enhancements include adjusting some key hyperparameters, eliminating the NSP objective, and training with a larger batch size and learning rate. ALBERT [31] introduces two parameter reduction techniques to reduce memory consumption and enhance the training speed of BERT.
@@
Encoder-Decoder-based PLMs: This neural architecture is primarily designed for sequence-to-sequence tasks, including machine translation, text summarization, and dialogue generation. These models integrate both the encoder and decoder modules of the Transformer, where the encoder processes the input sequence into continuous representations that capture contextual information, and the decoder sequentially generates the output sequence based on these representations. T5 [52] and BART [35] are two prominent Encoder-Decoder-based PLMs that have demonstrated exceptional performance in sequence-to-sequence tasks.
@@
The Text-to-Text Transfer Transformer (T5) model [52] advances the field of transfer learning in NLP by proposing a unified framework that reformulates all text-based language tasks into a text-to-text format. BART [35] utilizes a standard sequence-to-sequence model architecture augmented with a denoising strategy, in which the input text is intentionally corrupted using various noising functions such as token masking, document rotation, or sentence permutation. The model is then trained to reconstruct the original text from the corrupted input.
@@
Decoder-based PLMs: In these models, the attention layers at each stage are restricted to attending only to preceding words in the sentence, characterizing them as unidirectional or auto-regressive models. The pre-training process generally involves predicting the next word (or token) in the sequence. Consequently, decoder-based models are particularly effective for text generation tasks. The GPT [50, 51, 7] and LLaMA [65, 66] families have developed several powerful foundational models that utilize the Transformer’s decoder architecture. These models are pre-trained on extensive datasets comprising trillions of tokens and enhance the architecture through various techniques, such as employing the SwiGLU activation function instead of ReLU, incorporating rotary positional embeddings in place of absolute positional embeddings, and utilizing root-mean-squared layer normalization instead of the standard layer normalization.
@@
Large Language Models (LLMs) primarily refer to Transformer-based PLMs characterized by their extensive architecture, containing billions of parameters. These models are primarily inspired by decoder-based architectures, forming the foundation for the development of more advanced LLMs. LLMs are considerably larger in size, exhibiting superior language understanding and generation capabilities compared to small-scale PLMs. Some notable LLMs include GPT-4 [1], LLaMA-2 [66], PaLM [11], and FLAN [74].
@@
NL-based PLMs: This is a class of models primarily trained on extensive natural language text corpora [17, 39, 31, 52, 35, 50, 51, 7, 65, 66]. These models leverage vast amounts of textual data to learn rich linguistic representations, making them highly effective for a wide range of NLP tasks, such as text classification, sentiment analysis, and question answering.
Figure 1: An Overview of the Tesoro Creation Pipeline.
Code-based PLMs: These are specialized models designed to understand and generate programming code [22, 23, 24, 46, 70, 71, 54, 40, 75]. Typically, these models are initialized from NL-based PLMs and further trained on large corpora of source code from various programming languages. The datasets are collected from rich code sources, including GitHub and Stack Overflow. These models demonstrate exceptional performance across various code-related tasks, including code summarization, code translation, bug detection, technical debt detection, and code generation.
3 Proposed Methodology

### v2

@@ Abstract.
Technical debt (TD) describes the additional costs that emerge when developers have opted for a quick and easy solution to a problem, rather than a more effective and well-designed, but time-consuming approach. Self-Admitted Technical Debts (SATDs) are a specific type of technical debts that developers intentionally document and acknowledge, typically via textual comments. While these comments are a useful tool for identifying TD, most of the existing approaches focus on capturing tokens associated with various categories of TD, neglecting the rich information embedded within the source code. Recent research has focused on detecting SATDs by analyzing comments, and there has been little work dealing with TD contained in the source code. In this study, through the analysis of comments and their source code from 974 Java projects, we curated the first ever dataset of TD identified by code comments, coupled with its code. We found that including the classified code significantly improves the accuracy in predicting various types of technical debt. We believe that our dataset will catalyze future work in the domain, inspiring various research related to the recognition of technical debt; The proposed classifiers may serve as baselines for studies on the detection of TD.
Technical Debt, Pre-trained Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ; ; $c$s: Software and its engineering Software verification and validationccs: Software and its engineering Software testing and debugging
1. Introduction
@@
The concept of technical debt was originally introduced by Cunningham (Cunningham, 1992) to represent the liabilities that arise when developers make sub-optimal technical decisions, either intentionally or unintentionally during the software development life-cycle in their rush to market. Various factors can lead to the accumulation of technical debt, including deadline pressures, existing low-quality code, misaligned incentives, and poor software processes, among others (Bellomo et al., 2016). Previous studies have shown that developers often underestimate the consequences of such debts, which can degrade the quality of the source code, increase bug rates, and slow development velocity (Zazworka et al., 2011; Wehaibi et al., 2016). Identifying the code that contains technical debt is crucial to a rational development process, as this allows developers to fix the most important issues, the ones that are slowing the project down.
@@
In 2015, da Silva Maldonado et al. (Maldonado and Shihab, 2015) introduced their seminal work on detecting self-admitted technical debt (SATD) from comments embedded in source code. In particular, the authors manually classified a set of code comments to identify various types of technical debt. Afterward, they also developed various models to detect SATD from their dataset  (da Silva Maldonado et al., 2017). Significant research on the recognition of SATD has flourished since then (Di Salle et al., 2022; Sala et al., 2021). Recently, there have been various approaches proposed to recognize technical debt contained in different project artifacts. Among others, Li et al. (Li et al., 2023) conceived an approach to identify SATD from four independent sources, i.e., source code comments, commit messages, pull requests, and issue tracking systems. Tan et al. (Tan et al., 2023) manually curated a dataset of TD manifested in 3,000 issues. An evaluation of the collected dataset showed that there is a positive correlation between the number of TD items identified and mentioned as resolved in issue trackers and the number of debt items paid back in the source code.
@@
The majority of research thus far has mined TD from textual sources, e.g., comments (Maldonado and Shihab, 2015), issues (Tan et al., 2023; Li et al., 2023), or pull requests (Li et al., 2023). In this respect, SATD detection tools rely heavily on text to function. If no technical debt is reported in text, a debt might exist, but these tools would fail to identify it. To add insult to injury, even when there are comments, many of them and the corresponding code are not coherent, i.e., comments may be out of date and incorrectly reflect what is actually contained in the associated code (Corazza et al., 2018). This may happen because developers forget to update their comments after they made changes to the code (Igbomezie et al., 2024; Rabbi and Siddik, 2020).
There have been various attempts to associate SATD with weaknesses. Russo et al. (Russo et al., 2022) conceived WeakSATD to analyze source code written in C contained in Chromium projects to understand whether code blocks associated with SATD comments may contain weaknesses. The authors curated a set of heuristics from the public Common Weaknesses and Enumeration (CWE) repository to detect known weaknesses in software code and recommend mitigations. Other authors have investigated the possibility of detecting technical debt directly in source code. Nevertheless, these studies typically focused on limited classification schemes–such as distinguishing only between high and not high TD (Tsoukalas et al., 2024, 2022), or identifying the mere presence or absence of code smells (Yadav et al., 2024), without considering the broader diversity of technical debt types. To the best of our knowledge, no work has been conducted to create large-scale datasets of technical debt directly contained in Java source code. But we see a need for this type of data, to reduce the dependence on textual comments in detecting TD, thus vastly enhancing the contexts in which debt may be detected.
@@
To address these challenges, we conceive a new way of detecting technical debt. In particular, we propose a pipeline for the enrichment of technical debt data. Our methodology involves extracting SATD comments in conjunction with corresponding source code units. We have devised a method to identify Java source code that possibly contains technical debt, creating our initial corpus. Then we manually classified five categories of technical debt in this corpus. In addition, we developed a machine learning based tool to detect technical debt contained in textual comments and source code. By means of an empirical evaluation, we demonstrated that the curated dataset has the potential to advance state-of-the-art research in the domain, paving the way for a completely new way of identifying TD.
@@
  • RQ1: Does the inclusion of source code help to enhance the detection of technical debt? We enriched the input data with source code and fine-tuned four machine learning models, i.e., BERT, RoBERTa, UniXCoder, CodeBERT to identify technical debt in code. This RQ aims to investigate whether the enriched dataset (consisting of classified comments and corresponding source code) is beneficial to the detection of technical debt.
  • RQ2: What is the accuracy of different pre-trained models when detecting TD solely from source code? Using various machine learning models, we ran experiments on the collected dataset to investigate the accuracy of these models in classifying debt contained solely in source code. With this RQ we investigate to what extent existing deep learning algorithms are able to detect TD from code, thus inspiring future research in this direction.
  • RQ3: How do the manually classified comments contribute to the detection of SATD? We augmented the dataset collected by da Silva Maldonado et al. (Maldonado and Shihab, 2015) with our newly classified comments to yield a combined dataset. Afterward, we ran four machine learning models, i.e., BERT, RoBERTa, UniXCoder, CodeBERT, on both datasets to compare the prediction performance of these models. This aims to determine whether the new comments are useful in predicting technical debt.
@@ 2. Background
2.1. Self-admitted technical debt (SATD)
SATD is the technical debt that is expressly admitted by a developer through comments embedded in source code, issue trackers (Li et al., 2020), commit messages, or pull requests (Li et al., 2023). da Silva Maldonado et al. (Maldonado and Shihab, 2015) identified five types of SATD, i.e., DESIGN, DEFECT, DOCUMENTATION, REQUIREMENT/IMPLEMENTATION, and TESTING. This categorization allows for more insightful descriptions and a deeper understanding of the non-optimal solution options taken. This section takes various examples to explain these SATD categories.
@@ 2.1. Self-admitted technical debt (SATD)
2.2. Pretrained Language Models
Language models (LMs) are a foundational component in natural language processing (NLP) that have significantly advanced over the past decade. Recently, LMs have been powered by neural networks and trained on large text corpora, being able to capture both the syntactic and semantic aspects of languages more effectively. These models commonly follow the pre-training and fine-tuning paradigm (Zhao et al., 2023). During the pre-training phase, models are trained on large-scale unlabeled corpora using task-agnostic objectives such as word prediction, resulting in the development of pre-trained language models (PLMs). PLMs are then fine-tuned to adapt to various downstream tasks. Early PLMs (Sarzynska-Wawer et al., 2021) were mostly based on Recurrent Neural Networks (RNNs) and their variants, such as long short-term memory (LSTM) (Le Quoc et al., 2014) and gated recurrent units (GRU) (Cho et al., 2014). However, these approaches were computationally inefficient due to limitations in parallel processing, reducing scalability when training with extensive datasets and large model sizes. With the introduction of the Transformer architecture (Vaswani, 2017) and its self-attention mechanism, significantly more parallelization became possible as compared to RNNs. This advancement enables efficient pre-training of large language models on extensive datasets using multiple GPUs. Various transformer-based PLMs have achieved state-of-the-art performance across a wide range of tasks (Devlin et al., 2018; Liu et al., 2019; Lan et al., 2019; Raffel et al., 2020; Touvron et al., 2023a). Given the superior performance of transformer-based PLMs, which have also been explored in the context of SATD detection (Di Salle et al., 2022; Sharma et al., 2022; Sheikhaei et al., 2024), our study concentrates on utilizing these models.
2.2.1. Transformer architecture
This section provides an overview of the Transformer architecture, emphasizing key components and elements (Vaswani, 2017).
@@ 2.2.2. Types of PLMs
Based on the neural architectures of Transformer-based PLMs, we categorize the models into three main groups, as also outlined in existing work (Minaee et al., 2024).
@@
Encoder-based PLMs: This type of model utilizes the Transformer Encoder and builds a network by stacking multiple layers. These models were initially developed for language understanding tasks, such as text classification, where the objective is to predict a class label for a given input text. The pre-training stage of these models typically involves corrupting a given sentence in some way (e.g., by masking random words) and then training the model to identify or reconstruct the original sentence. To tackle a downstream task such as sentence classification or named entity recognition, these models are fine-tuned on task-specific data, and this involves substituting the LM head (the word prediction layer), with a classification head. BERT (Devlin et al., 2018), a prominent encoder-based model, has inspired the development of several variants, such as RoBERTa (Liu et al., 2019) and ALBERT (Lan et al., 2019), which have demonstrated substantial improvements across various understanding tasks.
@@
Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2018) is among the most widely adopted encoder-based PLMs. During pretraining, BERT leverages two objectives: masked language modeling (MLM) and next sentence prediction (NSP). In MLM, random tokens within a sentence are masked, and the model is trained to predict these masked tokens using the context of the surrounding words. Meanwhile, NSP trains BERT to comprehend the relationship between two sentences by predicting whether one sentence logically follows the other. RoBERTa (Liu et al., 2019) extends BERT by improving its robustness through refined model design choices and training strategies. These enhancements include adjusting some key hyperparameters, eliminating the NSP objective, and training with a larger batch size and learning rate. ALBERT (Lan et al., 2019) introduces two parameter reduction techniques to reduce memory consumption and enhance the training speed of BERT.
@@
Encoder-Decoder-based PLMs: This neural architecture is primarily designed for sequence-to-sequence tasks, including machine translation, text summarization, and dialogue generation. These models integrate both the encoder and decoder modules of the Transformer, where the encoder processes the input sequence into continuous representations that capture contextual information, and the decoder sequentially generates the output sequence based on these representations. T5 (Raffel et al., 2020) and BART (Lewis, 2019) are two prominent Encoder-Decoder-based PLMs that have demonstrated exceptional performance in sequence-to-sequence tasks.
@@
The Text-to-Text Transfer Transformer (T5) model (Raffel et al., 2020) advances the field of transfer learning in NLP by proposing a unified framework that reformulates all text-based language tasks into a text-to-text format. BART (Lewis, 2019) utilizes a standard sequence-to-sequence model architecture augmented with a denoising strategy, in which the input text is intentionally corrupted using various noising functions such as token masking, document rotation, or sentence permutation. The model is then trained to reconstruct the original text from the corrupted input.
@@
Decoder-based PLMs: In these models, the attention layers at each stage are restricted to attending only to preceding words in the sentence, characterizing them as unidirectional or auto-regressive models. The pre-training process generally involves predicting the next word (or token) in the sequence. Consequently, decoder-based models are particularly effective for text generation tasks. The GPT (Radford, 2018; Radford et al., 2019; Brown, 2020) and LLaMA (Touvron et al., 2023a, b) families have developed several powerful foundational models that utilize the Transformer’s decoder architecture. These models are pre-trained on extensive datasets comprising trillions of tokens and enhance the architecture through various techniques, such as employing the SwiGLU activation function instead of ReLU, incorporating rotary positional embeddings in place of absolute positional embeddings, and utilizing root-mean-squared layer normalization instead of the standard layer normalization.
@@
Large Language Models (LLMs) primarily refer to Transformer-based PLMs characterized by their extensive architecture, containing billions of parameters. These models are primarily inspired by decoder-based architectures, forming the foundation for the development of more advanced LLMs. LLMs are considerably larger in size, exhibiting superior language understanding and generation capabilities compared to small-scale PLMs. Some notable LLMs include GPT-4 (Achiam et al., 2023), LLaMA-2 (Touvron et al., 2023b), PaLM (Chowdhery et al., 2023), and FLAN (Wei et al., 2021).
@@
Figure 1. An Overview of the Tesoro Creation Pipeline.
  • NL-based PLMs: This is a class of models primarily trained on extensive natural language text corpora (Devlin et al., 2018; Liu et al., 2019; Lan et al., 2019; Raffel et al., 2020; Lewis, 2019; Radford, 2018; Radford et al., 2019; Brown, 2020; Touvron et al., 2023a, b). These models leverage vast amounts of textual data to learn rich linguistic representations, making them highly effective for a wide range of NLP tasks, such as text classification, sentiment analysis, and question answering.
  • Code-based PLMs: These are specialized models designed to understand and generate programming code (Feng et al., 2020; Guo et al., 2022, 2020; Nijkamp et al., 2022; Wang et al., 2023, 2021; Roziere et al., 2023; Lozhkov et al., 2024; Wei et al., 2023). Typically, these models are initialized from NL-based PLMs and further trained on large corpora of source code from various programming languages. The datasets are collected from rich code sources, including GitHub and Stack Overflow. These models demonstrate exceptional performance across various code-related tasks, including code summarization, code translation, bug detection, technical debt detection, and code generation.
3. Proposed Methodology